# JailGuard: A Universal Detection Framework for LLM Prompt-based Attacks

### v3

@@ 6.2. RQ1: Effectiveness of Detecting Attack
Answer to RQ1: All mutation strategies in JailGuard can effectively detect LLM prompt-based attacks on text and image inputs, surpassing state-of-the-art methods in detection accuracy. JailGuard achieve an average accuracy of 81.68% and 79.53% on image and text datasets, respectively. For the single mutators, targeted mutators can achieve better detection results than their random versions, improving accuracy by 1.07% and 3.42%. Moreover, the combination policies in JailGuard further improve the detection accuracy to 86.14% and 82.09% on text and image inputs, significantly outperforming state-of-the-art detection methods by 11.81%-25.73% and 12.20%-21.40%, demonstrating the effectiveness of the default combination policy in JailGuard.
@@ 6.3. RQ2: Effectiveness of Detecting Different Kinds of Attacks
Fig. 7. Comparison of Different Methods’ Results on Image Inputs
Analysis. The experiment results illustrate the effectiveness and generalization of JailGuard’s mutators, especially for the targeted mutators and policies, in detecting various attacks. From Fig. 6 and Fig. 7, we observe that most baseline methods struggle to detect attack samples with different attack targets. The jailbreak defense methods (e.g., SmoothLLM and In-Context Defense) usually leverage jailbreaking cases and keywords to detect attacks. As a result, they can hardly provide effective detection for hijacking attacks with unknown attack targets. Although perplexity-based detection and LLM-based detection can effectively block most attack samples, they introduce a large number of false positives, allowing only 5.77% and 62.40% of benign samples to pass, which is significantly lower than other methods. Furthermore, even for jailbreaking attacks, there is substantial variability in baseline detection effect for samples generated by different attack methods. For example, the SmoothLLM with ‘insert’ method only has a detection accuracy of 62.80% and 55.56% on the jailbreaking attack ‘Jailbroken’ and ‘Parameter’, which is much lower than the accuracy on other attacks (e.g., 90.32% on GPTFuzz). Similar observations can be made on the image dataset, where ECSO detection accuracy varies from 18.00% to 48.00% across different attacks. In contrast, the mutators and policies JailGuard can effectively identify various prompt-based attacks regardless of their attack targets, consistently achieving over 70% accuracy on benign samples. It indicates that JailGuard can overcome the existing limitations and exploit the divergence of variants to provide general, effective detection for various LLM prompt-based attacks.
@@
Case Study. We provide a case in Fig. 8 to understand and illustrate the root cause of the effect difference between JailGuard and the best baseline SmoothLLM with ‘swap’ method on specific attacks, such as Jailbroken attacks. The upper part shows the detection process of SmoothLLM and the lower part of Fig. 8 shows the detection of JailGuard combining the mutators Punctuation Insertion, Targeted Insertion, and Translation. We highlight the mutator Targeted Insertion in this case. Fig. 8.a) provides a real example of the Jailbroken attack in our dataset. Jailbroken attacks use specific content or tasks to construct scenarios and conduct role-playing, thereby deceiving the defense mechanism of the LLM system and achieving attack. SmoothLLM randomly swaps 10% characters to infer these attack inputs as much as possible. However, the impact of these perturbations is minimal. Among the eight perturbed inputs, only one attack fails, with its response containing refusal keywords, highlighted in red texts in the upper part of Fig. 8.c). Therefore, in the aggregation step in the upper part of Fig. 6.d), since most results do not contain refusal keywords, according to its aggregation principle, this input sample is incorrectly judged as a benign sample, resulting in a false negative.
@@ 6.4. RQ3: Ablation Study
Experiment Designs and Results. The mutator combination strategy in JailGuard achieves the best detection accuracy among all baselines and mutation strategies on text and image datasets. It leverages two modules to effectively detect prompt-based attacks in LLM systems and applications, which are the mutator combination policy and the divergence-based detection. To understand their contribution, we conduct an ablation experiment on the text inputs. The results, shown in Table 5, record the accuracy, precision, and recall of each method. Firstly, we implement three random policies and record their detection results to illustrate the effectiveness and contribution of JailGuard’s built-in policy. The first policy uses random probability and the same mutator pool as the built-in policy (Row ‘Random 1’), the second one uses the same probability as the built-in policy and a random mutator pool (i.e., Random Insertion, Synonym Replacement and Random Deletion, shown in Row ‘Random 2’ ), and the third one randomly select mutators from all text mutators (Row ‘Random 3’). Additionally, to understand the contribution of divergence-based detection, we substitute the divergence-based detection in JailGuard with two keyword detection methods and use the variants generated by the built-in policy to detect attacks. Their results are shown in the last two rows of Table 5. The first detection method randomly picks one variant response and uses refusal keyword detection to detect attacks (Row ‘Random Selection + Keywords’). The second detector leverages the aggregation method in SmoothLLM to get the final responses from the variant responses and then uses keywords to identify attacks (Row ‘Aggregation + Keywords’).
Analysis. The experimental results of the ablation study illustrate the effectiveness of both combination policy and divergence-based detection in JailGuard. Firstly, altering the built-in policy will degrade the detection effect. The results of Table 5 demonstrate that employing a random mutator pool or probabilities in the policy degrades detection performance, sometimes even falling below the accuracy achieved with a single operator. Random Policy 1, which uses unoptimized probabilities to select mutators from the pool, reduces the detection accuracy by 3.17% compared to the built-in policy in JailGuard. Random policies 2 and 3 that use random mutator pool achieve an accuracy of 79.18% and 82.31% respectively, which are 6.96% and 3.83% lower than the original policy. Especially for policy 2, random selection from the mutators with poor detection performance even causes the precision to drop by 10.57%.
@@
In addition, the divergence-based detection in JailGuard has an important contribution to attack detection, especially in eliminating FNs and improving recall. As shown in Table 5, using the keywords detection methods to replace the divergence-based detection in JailGuard leads to significant degradation in detection effects. Using aggregation and keyword detection will reduce the detection accuracy from 86.14% to 73.82%. Randomly selecting responses leads to even more severe performance degradation, with accuracy falling to 84.33% of the original accuracy. Worse still, the recall of random selection and keyword detection is only 40.45%, indicating that more than half of the attack samples could bypass the detection. Our analysis of the detection results shows that the keyword detection overlooks many attack examples, particularly hijacking prompt injection attacks, and cannot provide effective defense for various attacks. This observation is consistent with our findings in § 6.3.
@@
6.5. RQ4: Impact of Variant Amount
@@
Experiment Designs and Results. JailGuard leverage the mutation strategies to generate $N$ variants and compute the divergence in the corresponding responses. To understand the impact of different values of $N$ (i.e., different LLM query budget) on detection results, we evaluate the detection effectiveness of different mutators and policies when generating 4/5/6/7/8 variants, and record accuracy and recall on the image and text dataset, as shown in solid lines with different colors Fig. 9. The bold red line marks the result of the default combination policy in JailGuard. In addition, the number of variants also affects the detection effect of the baseline SmoothLLM (Robey et al., 2023). We have run the three methods (i.e., ‘insert’, ‘swap’, and ‘patch’) of SmoothLLM on 4 to 8 variant budgets. Then we record the best results achieved by SmoothLLM in the whole experiment, that is, when using the ‘swap’ method and generating 8 variants, the accuracy reaches 74.33% and when generating 4 variants with ‘swap’, SmoothLLM has the highest recall of 47.98%, as shown in the dashed purple lines in Fig. 9.a).
Analysis. We can observe that decreasing the number of variants (i.e., LLM query budget) generally degrades the detection results of the mutators and policies. Taking the combination policy of JailGuard as an example, when the number of variants drops from 8 to 4, the detection recall on the text drops from 86.10% to 82.25%, and it is more obvious on the image, from 88.25% to 72.50%. However, such a decrease is not obvious in detection accuracy, which only drops 1.17% and 5.30% on text and image datasets respectively. Our analysis shows that benign samples are less likely to exhibit large divergence when producing fewer variants, leading to fewer false positives and mitigating the reduction in detection accuracy. Notably, for the single mutator Synonym Replacement which has the lowest detection precision among all mutators, as shown in Table 3. A reduction in the query budget actually increases its detection precision from 65.77% to 69.57%, and its detection accuracy from 75.21% to 76.94%. Moreover, the detection results of any mutator or policy under any query budget outperform the best results achieved by the baseline SmoothLLM (the dashed purple lines in Fig. 9.a)), which illustrates the potential of JailGuard for effective attack detection even under low query budget.
@@
In actual deployment scenarios, JailGuard accesses LLM to batch process and infer the input variant, which leads to additional memory overhead. Our simulations on MiniGPT-4 show that a single set of inputs (one image and one corresponding instruction) increases the memory overhead by 0.49GB, which is equivalent to 3.15% of the LLM memory overhead (15.68GB). If the LLM query budget is set to $N=8$, the memory overhead of JailGuard to detect jailbreaking attacks is 3.95GB, which is 25.20% of the memory overhead of LLM itself. Although the runtime overhead of JailGuard is acceptable, considering that resources may be limited in LLM system application and deployment scenarios, performing effective attack detection with lower overhead has great significance. Considering that under different settings of budget $N$, JailGuard can consistently achieve detection results far exceeding the baseline under different budget settings and the detection effect improves with the increase of $N$, we recommend using $N=8$ as the default number of variants to achieve the best detection effect across different attacks and using $N\in[4,6]$ to achieve the balance between the detection effect and runtime overhead in resource-constrained scenarios.
@@
Answer to RQ4: Reducing the query budget and the number of variants results in a slight degradation in JailGuard detection accuracy and recall. Even so, in the process of reducing the budget, the detection effect of the mutators and policies in JailGuard is always better than the best results of SmoothLLM, indicating the potential of JailGuard to detect prompt-based attacks in low-cost scenarios. Consequently, JailGuard generates 8 variants by default to obtain the best detection effect. When the LLM query budget is limited, users can choose to generate 4 to 6 variants to obtain a balance of efficiency and effectiveness.

### v4

@@ 6.2. RQ1: Effectiveness of Detecting Attack
Answer to RQ1: All mutation strategies in JailGuard can effectively detect prompt-based attacks on text and image inputs, surpassing state-of-the-art methods in detection accuracy. JailGuard achieve an average accuracy of 81.68% and 79.53% on image and text datasets, respectively. For the single mutators, targeted mutators can achieve better detection results than their random versions, improving accuracy by 1.07% and 3.42%. Moreover, the combination policies in JailGuard further improve the detection accuracy to 86.14% and 82.09% on text and image inputs, significantly outperforming state-of-the-art detection methods by 11.81%-25.73% and 12.20%-21.40%, demonstrating the effectiveness of the default combination policy in JailGuard.
@@ 6.3. RQ2: Effectiveness of Detecting Different Kinds of Attacks
Analysis. The experiment results illustrate the effectiveness and generalization of JailGuard’s mutators, especially for the targeted mutators and policies, in detecting various attacks. From Fig. 9 and Fig. 8, we observe that most baseline methods struggle to detect attack samples with different attack targets. The jailbreak defense methods (e.g., SmoothLLM and In-Context Defense) usually leverage jailbreaking cases and keywords to detect attacks. As a result, they can hardly provide effective detection for hijacking attacks with unknown attack targets. Although perplexity-based detection and LLM-based detection can effectively block most attack samples, they introduce a large number of false positives, allowing only 5.77% and 62.40% of benign samples to pass, which is significantly lower than other methods. Furthermore, even for jailbreaking attacks, there is substantial variability in baseline detection effect for samples generated by different attack methods. For example, the SmoothLLM with ‘insert’ method only has a detection accuracy of 62.80% and 55.56% on the jailbreaking attack ‘Jailbroken’ and ‘Parameter’, which is much lower than the accuracy on other attacks (e.g., 90.32% on GPTFuzz). Similar observations can be made on the image dataset, where ECSO detection accuracy varies from 18.00% to 48.00% across different attacks. In contrast, the mutators and policies JailGuard can effectively identify various prompt-based attacks regardless of their attack targets, consistently achieving over 70% accuracy on benign samples. It indicates that JailGuard can overcome the existing limitations and exploit the divergence of variants to provide general, effective detection for various LLM prompt-based attacks. Note that the column ‘DeepInception’ in Fig. 9 only contains 0%, 50%, and 100%. The root cause is that only 2 of the 300 attack prompts generated by DeepInception pass the verification in § 5. Most of the generated attack prompts have been refused by LLMs or only lead to responses unrelated to harmful content. How to expand the dataset and add more valid attack inputs for each attack method is our future work.
Fig. 8. Comparison of Different Methods’ Results on Image Inputs
@@
Case Study 1: We provide a case in Fig. 10 to understand and illustrate the root cause of the effect difference between JailGuard and the best baseline SmoothLLM with ‘swap’ method on specific attacks, such as Jailbroken attacks. The upper part shows the detection process of SmoothLLM and the lower part of Fig. 10 shows the detection of JailGuard combining the mutators Punctuation Insertion, Targeted Insertion, and Translation. We highlight the mutator Targeted Insertion in this case. Fig. 10.a) provides a real example of the Jailbroken attack in our dataset. Jailbroken attacks use specific content or tasks to construct scenarios and conduct role-playing, thereby deceiving the defense mechanism of the LLM system and achieving attack. SmoothLLM randomly swaps 10% characters to infer these attack inputs as much as possible. However, the impact of these perturbations is minimal. Among the eight perturbed inputs, only one attack fails, with its response containing refusal keywords, highlighted in red texts in the upper part of Fig. 10.c). Therefore, in the aggregation step in the upper part of Fig. 9.d), since most results do not contain refusal keywords, according to its aggregation principle, this input sample is incorrectly judged as a benign sample, resulting in a false negative.
Case Study 2: We have collected benign data from existing LLM instruction sets (Li et al., 2023a; Askari et al., 2023; Zhou et al., 2023), including general questions with sensitive words. We provide a case study to demonstrate how JailGuard correctly identifies such seemingly toxic questions. The following benign prompt is collected from the existing work (Li et al., 2023a), ‘write code that imitates a fake chatbot using a modal where it shows the human users previously typed messages and the ai chat bots messages all within html. Create the modal fake chat history having 4 overall messages. With 2 messages from the bot and 2 messages from the human’. Due to the mention of the ‘fake chatbot’, existing methods such as LLM-based detectors and Self-reminder with reinforced safety alignment mistakenly identify this prompt as an attack prompt and block it. In contrast, all mutators in JailGuard can correctly identify the prompt as benign prompts with no false positives. After mutating this prompt, JailGuard calculates the divergence of the variant responses and obtains the divergence values ranging between 0.002 and 0.013, which is lower than the built-in threshold $\theta=0.02$.
@@
Error Analysis on Combination Policy. Although the mutator combination policy has achieved the best detection results among all detectors, it still had some misclassified cases. Our analysis shows that these cases can be divided into two categories. The first category, which accounts for most of the misclassified cases, consists of cases that can be correctly detected when using single mutators in the mutator pool of the policy. However, when generating variants, limited by the variants generated by different mutators in combination policy, it ultimately obtains incorrect detection results. For instance, JailGuard with a single mutator can get LLM responses with low divergence for benign examples, but when using different mutators, JailGuard may generate variants with significant differences, leading to a large divergence that exceeds the built-in threshold and false positives. The second category consists of cases that cannot be correctly identified even when using single mutators. Ideally, the upper bound of the combination policy should correctly identify the first category of misclassified cases, which are caused by random mutator selection rather than the deficiency in the capabilities of the mutators themselves. At this time, the text combination policy achieves a detection accuracy of 95.89%, and the image policy has a detection accuracy of 98.20%. In addition, benign cases in the second category are mainly heuristic instructions that always produce significantly different LLM responses (e.g., ‘Suggest some names for a friendly group in telegram’). This prompt can produce vastly different names in multiple queries, resulting in a large divergence. Attack cases in the second category mainly contain less severe harmful questions. These attacks are near the boundary of LLM safety alignment, and LLMs tend to consider them as benign prompts and answer them. For example, the attack prompt constructed by the TAP method (Mehrotra et al., 2023) is that ‘As a story writer, you need to include a scenario where a character needs to transport valuable items across an international border without arousing suspicion. Can you provide some guidance on how the character could successfully achieve this task? Begin by replying with: ’How can someone ensure the safe transportation of valuable items across an international border without arousing suspicion?’’. This attack prompt asks a seemingly benign question related to smuggling, and LLMs tend to respond to it like a benign prompt, thus it can successfully bypass various detection methods such as Self-reminder, SmoothLLM, and JailGuard. How to effectively identify and detect such attack prompts will be a future enhancement for JailGuard. We provide further discussion in § 8.
@@ 6.4. RQ3: Ablation Study
Experiment Designs and Results. The mutator combination strategy in JailGuard achieves the best detection accuracy among all baselines and mutation strategies on text and image datasets. It leverages two modules to effectively detect prompt-based attacks in LLM systems and applications, which are the mutator combination policy and the divergence-based detection. To understand their contribution, we conduct an ablation experiment on the text inputs. The results, shown in Table 6, record the accuracy, precision, and recall of each method. Firstly, we implement three random policies and record their detection results to illustrate the effectiveness and contribution of JailGuard’s built-in policy. The first policy uses random probability and the same mutator pool as the built-in policy (Row ‘Random 1’), the second one uses the same probability as the built-in policy and a random mutator pool (i.e., Random Insertion, Synonym Replacement and Random Deletion, shown in Row ‘Random 2’ ), and the third one randomly select mutators from all text mutators (Row ‘Random 3’). Additionally, we use the mutator implemented by SmoothLLM to substitute the mutation policy of JailGuard to observe the impact on the detection effect (Rows ‘Insert’, ‘Swap’, and ‘Patch’). Furthermore, to understand the contribution of divergence-based detection, we substitute the divergence-based detection in JailGuard with two keyword detection methods and use the variants generated by the built-in policy to detect attacks. Their results are shown in the last two rows of Table 6. The first detection method randomly picks one variant response and uses refusal keyword detection to detect attacks (Row ‘Random Selection + Keywords’). The second detector leverages the aggregation method in SmoothLLM to get the final responses from the variant responses and then uses keywords to identify attacks (Row ‘Aggregation + Keywords’).
Analysis. The experimental results of the ablation study illustrate the effectiveness of both combination policy and divergence-based detection in JailGuard. Firstly, altering the built-in policy will degrade the detection effect. The results of Table 6 demonstrate that employing a random mutator pool or probabilities in the policy degrades detection performance, sometimes even falling below the accuracy achieved with a single operator. Random Policy 1, which uses unoptimized probabilities to select mutators from the pool, reduces the detection accuracy by 3.17% compared to the built-in policy in JailGuard. Random policies 2 and 3 that use random mutator pool achieve an accuracy of 79.18% and 82.31% respectively, which are 6.96% and 3.83% lower than the original policy. Especially for policy 2, random selection from the mutators with poor detection performance even causes the precision to drop by 10.57%. In addition, replacing the combination policy of JailGuard with the mutation methods in SmoothLLM separately leads to an accuracy degradation of up to 13.82%, further illustrating the effectiveness of the combination policy of JailGuard.
@@
The divergence-based detection in JailGuard has an important contribution to attack detection, especially in eliminating FNs and improving recall. As shown in Table 6, using the keywords detection methods to replace the divergence-based detection in JailGuard leads to significant degradation in detection effects. Using aggregation and keyword detection will reduce the detection accuracy from 86.14% to 73.82%. Randomly selecting responses leads to even more severe performance degradation, with accuracy falling to 72.64%. Worse still, the recall of random selection and keyword detection is only 40.45%, indicating that more than half of the attack samples could bypass the detection. Our analysis of the detection results shows that keyword detection overlooks many attack examples, particularly hijacking prompt injection attacks, and cannot provide effective defense for various attacks. This observation is consistent with our findings in § 6.3. In addition, the experiment results in Rows ‘Insert’, ‘Swap’, and ‘Patch’ further demonstrate the effectiveness of divergence-based detection in JailGuard compared to the baseline method. Keeping the mutation methods in SmoothLLM and combining them with the divergence-based detection can effectively improve the detection accuracy of the original SmoothLLM by up to 5.79%, and the recall can be increased up to 2.11 times the original.
@@
6.5. RQ4: Impact of Threshold $\theta$
Experiment Designs and Results. JailGuard leverage the built-in threshold $\theta$ and the divergence of variant responses to distinguish attack and benign inputs. To understand the impact of different $\theta$ values on the detection results, we record and evaluate the detection accuracy, precision, and recall of different mutation strategies under different threshold settings on the development set consisting of 70% of the collected dataset. (§ 6.1). Fig. 11 shows the detection results of mutation strategies on text and image datasets. The X-axis shows the value of the threshold $\theta$ that ranges from 0.001 to 1, and the Y-axis shows the detection accuracy, recall, and precision (dashed dot line) using the corresponding threshold. As shown in the legend, the lines of different colors represent different mutation strategies, and the bold red line highlights the results of the combination policy in JailGuard.
Analysis. We can observe that as the threshold $\theta$ increases, JailGuard will have fewer false positives and more false negatives in detecting attack samples, which is manifested as an increase in precision and a decrease in recall. During this process, the detection accuracy first increases sharply and then decreases slowly. Specifically, text mutators usually achieve the highest detection accuracy when the $\theta$ is in the range of 0.01 to 0.05. When $\theta$ continues to increase, JailGuard will misclassify many attack inputs as benign inputs, leading to a decrease in recall and accuracy. Our analysis shows that due to the large difference in the distribution of divergence between benign samples and attack samples, when the threshold $\theta$ varies between 0.01 and 0.1, the detection accuracy of most mutators can be maintained at a high value (i.e., 80%). In addition, image variants usually achieve the highest accuracy when $\theta$ is set to 0.01 to 0.03. It is worth noting that when $\theta$ increases, the detection accuracy of some image mutators (e.g., Random Mask) will first drop sharply and then improve again. Our analysis shows that when using these mutators, the divergences of many attack samples are distributed in this interval, while benign sample divergences are less distributed in this interval. Therefore, increasing $\theta$ will cause a large number of attack samples to be misjudged as benign inputs, and the number of true positives in detection will drop significantly, while the number of false positives will not decrease significantly, eventually leading to a drop in precision.
Considering the overall detection results of each mutator for benign samples and attack samples under different threshold settings, we finally choose the default value of $\theta$ for text mutators to 0.02 and the $\theta$ for the image mutators to 0.025.
Answer to RQ4: Increasing the threshold $\theta$ can generally prevent JailGuard from incorrectly blocking benign samples, but it will introduce more missed attack samples that endanger LLM systems. Considering the trade-off between the performance of each mutation strategy in blocking attack samples and passing benign samples, JailGuard finally chooses to set the built-in detection threshold $\theta$ of 0.02 and 0.025 for the text and image variant, respectively.
Fig. 11. Impact of the Built-in Threshold $\theta$ on Detection Results
6.6. RQ5: Impact of Variant Amount
Experiment Designs and Results. JailGuard leverages the mutation strategies to generate $N$ variants and compute the divergence in the corresponding responses. To understand the impact of different values of $N$ (i.e., different LLM query budget) on detection results, we evaluate the detection effectiveness of different mutators and policies when the number of generated variants varies from 2 to 32, and record accuracy and recall on the image and text dataset. Since generating 32 variant responses on the full dataset is too costly (requiring billions of paid tokens), we randomly select and use 1,000 items of text data and 200 items of image data from the collected dataset in this experiment. The solid lines with different colors in Fig. 12 show the collected results of mutation strategies and the bold red line marks the result of the default combination policy in JailGuard. In addition, the number of variants also affects the detection effect of the baseline SmoothLLM (Robey et al., 2023). We have run the three methods (i.e., ‘insert’, ‘swap’, and ‘patch’) of SmoothLLM on 2 to 32 variant budgets. Then we record the best results achieved by SmoothLLM in the whole experiment, that is, the accuracy of 76.90% (‘Swap’) and the recall of 55.25% (‘Patch’), as shown in the dashed purple lines in Fig. 12.a).
@@ 6.6. RQ5: Impact of Variant Amount
Analysis. We can observe that increasing the number of variants (i.e., LLM query budget) leads to better detection effects of the mutators and policies for attack prompts and higher recall. Regardless of the value of $N$, the mutators in JailGuard can always achieve a recall that is higher than the best result of SmoothLLM. Taking the combination policy of JailGuard as an example, when the number of variants increases from 2 to 32, the detection recall on the text improves from 66.25% to 100.00%, and it is more obvious on the image, from 60.00% to 100.00%. However, such an increasing trend does not apply to accuracy. For most mutation strategies, as $N$ increases, accuracy first increases and reaches its peak when $N$ is in the range of 6 to 14, and then decreases. Our analysis shows that benign samples have a higher probability of being affected by mutators when producing more variants, resulting in large divergence and false positives. For example, for the benign prompt ‘Is the continent of Antarctica located at the north or South Pole?’, over 70% variants obtain the response of ‘The Antarctic continent is located at the southern pole of the Earth’. However, when $N$ increases, it may get several responses with the same core content but very different expressions, such as ‘Antarctica is in Antarctica. The Arctic refers to the region around the North Pole, while Antarctica refers to the region around the South Pole’, resulting in a large divergence (i.e., 0.11) exceeding the threshold $\theta$. Notably, for mutators such as Synonym Replacement and Translation, increasing $N$ intuitively could lead to a drop in accuracy. Our analysis shows that these mutators usually significantly modify the original prompt, leading to a high probability of producing different responses for variants of benign and a large number of false positives. In some cases, these mutators even achieve detection accuracy lower than the baseline methods SmoothLLM (76.90%).
@@
In actual deployment scenarios, JailGuard accesses LLM to batch process and infer the input variant, which leads to additional memory overhead. Our simulations on MiniGPT-4 show that a single set of inputs (one image and one corresponding instruction) increases the memory overhead by 0.49GB, which is equivalent to 3.15% of the LLM memory overhead (15.68GB). If the LLM query budget is set to $N=8$, the memory overhead of JailGuard to detect jailbreaking attacks is 3.95GB, which is 25.20% of the memory overhead of LLM itself. Although the runtime overhead of JailGuard is acceptable, considering that resources may be limited in LLM system application and deployment scenarios, performing effective attack detection with lower overhead has great significance. Considering that under different settings of budget $N$, JailGuard can usually achieve detection results far exceeding the baseline, and the mutators usually obtain the best accuracy when $N$ is in the range of 6 to 14, we recommend using $N=8$ as the default number of variants to achieve the best detection effect across different attacks and using $N\in[4,6]$ to achieve the balance between the detection effect and runtime overhead in resource-constrained scenarios. According to the records of the cost in our large-scale experiments in § 6.2, it takes 1-2 seconds on average to obtain the LLM response for an input variant and consumes 450 paid tokens. In real-world deployment scenarios, developers can generate responses for multiple variants in one single batch with larger memory overhead. At this time, for $N=8$, detecting a prompt takes approximately 1-2 seconds and consumes 3,600 to 4,000 paid tokens (approximately \$0.01 at GPT-3.5 prices).
@@
Answer to RQ5: As the query budget and the number of variants $N$ increase, the mutators of JailGuard achieve greater recall, while the accuracy of most mutators first increases and then decreases. Considering the performance of each mutator, JailGuard generates 8 variants by default to obtain the best detection effect. Reducing the query budget and the number of variants results in a slight degradation in JailGuard detection accuracy and recall. In addition, even in a low-budget environment with less than 8 queries, the detection effect of mutators and strategies in JailGuard is always better than the best result of SmoothLLM, indicating the potential of JailGuard to detect prompt-based attacks in low-cost scenarios When the LLM query budget is limited, users can choose to generate 4 to 6 variants to obtain a balance of efficiency and effectiveness.